### v1

# Localized Deviations from the CO–PAH Relation in PHANGS–JWST Galaxies: Faint PAH Emission or Elevated CO Emissivity?

Abstract
Polycyclic aromatic hydrocarbon (PAH) emission is widely used to trace the distribution of molecular gas in the interstellar medium (ISM), exhibiting a tight correlation with CO(2–1) emission across nearby galaxies. Using PHANGS-JWST and PHANGS-ALMA data, we identify localized regions where this correlation fails, with CO flux exceeding that predicted from 7.7m PAH emission by more than an order of magnitude. These outlier regions are found in 20 out of 70 galaxies and are mostly located in galaxy centers and bars, without signs of massive star formation. We explore two scenarios to explain the elevated CO-to-PAH ratios, which can either be due to suppressed PAH emission or enhanced CO emissivity. We examine PAH emission in other bands (3.3m and 11.3m) and the dust continuum dominated bands (10m and 21m), finding consistently high CO-to-PAH (or CO-to-dust continuum) emission ratios, suggesting that 7.7 m PAH emission is not particularly suppressed. In some outlier regions, PAH size distributions and spectral energy distribution of the illuminating radiation differ slightly compared to nearby control regions with normal CO-to-PAH ratios, though without a consistent trend. We find that the outlier regions generally show higher CO velocity dispersions (). This increase in may lower CO optical depth and raise its emissivity for a given gas mass. Our results favor a scenario where shear along the bar lanes and shocks at the tips of the bar elevate CO emissivity, leading to the breakdown of the CO–PAH correlation.
show]jyeonkim@stanford.edu
]sutterjs@whitman.edu
1 Introduction
Emission from polycyclic aromatic hydrocarbons (PAHs) captured with JWST reveals a complex and widespread network of interstellar medium (ISM) structures throughout galaxy disks (J. C. Lee et al., 2023; D. A. Thilker et al., 2023; D. B. Fisher et al., 2024; T. G. Williams et al., 2024). In the diffuse Milky Way ISM, the fraction of cold neutral gas with respect to the total gas strongly correlates with PAH abundance, suggesting that PAHs are destroyed in warm gas while PAH growth and survival is facilitated in well-shielded, moderately dense gas (F. Galliano et al., 2018; B. S. Hensley & B. T. Draine, 2023a). In molecular gas-dominated galaxies, studies have shown a tight correspondence of PAH emission with CO emission as well as with H at scales ranging from a few kpc down to 100 pc (M. W. Regan et al., 2006; R. Chown et al., 2021; Y. Gao et al., 2022; A. K. Leroy et al., 2023; C. M. Whitcomb et al., 2023; R. Chown et al., 2025; H. Shim et al., 2025).
The tight correlation observed between CO and PAH emission in molecular gas-dominated galaxies can be understood within the standard framework of dust in galaxies, in which dust and PAH grains are well mixed within the gaseous phase of the ISM (B. T. Draine & A. Li, 2007; F. Galliano et al., 2018; B. S. Hensley & B. T. Draine, 2023a). PAHs are stochastically excited by ultraviolet and optical photons, allowing them to emit efficiently in the mid-infrared. As a result, the distribution of mid-IR emission is expected, to first order, to closely resemble the distribution of the cold ISM (A. K. Leroy et al., 2023; R. Chown et al., 2025). Indeed, PAH emission has even been used to probe the ISM in Hi-dominated regions with gas surface densities below (K. M. Sandstrom et al., 2023a).
The excitation of PAHs also depends on the intensity of the radiation field, leading to brighter PAH emission near star-forming regions (D. Pathak et al., 2024). A. K. Leroy et al. (2023) demonstrated that mid-IR emission can be described as a linear combination of scaled CO and H fluxes. However, strong ionizing radiation from young massive stars also destroys PAHs, as evidenced by suppressed PAH emission at 7.7 and 11.3m relative to the hot dust continuum at 21m in Hii regions (K. D. Gordon et al., 2008; J. Chastenet et al., 2019, 2023; O. V. Egorov et al., 2023, 2025; J. Sutter et al., 2024).
Recently, R. Chown et al. (2025) have characterized the relationship between 12CO(; hereafter CO) and PAH emission across 70 galaxies from the Physics at High Angular Resolution in Nearby Galaxies Survey (PHANGS). By combining Atacama Large Millimeter/sub-millimeter Array (ALMA) CO maps and JWST near- and mid-infrared (near-IR and mid-IR) imaging, R. Chown et al. (2025) have found a relatively small scatter () in the observed CO-to-PAH ratios at 100 pc-scale, where CO shows a strong and approximately linear correlation with 7.7m PAH emission, one of the common PAH tracers accessible with JWST. Similarly linear trends were observed at 11.3m and 3.3m for a subsample of 19 galaxies from the PHANGS–JWST Cycle 1 survey (J. C. Lee et al., 2023).
The growing evidence of PAH closely tracing CO has led to the suggestion that PAH emission can be used as a tracer of the cold ISM, not only for high–resolution studies of nearby galaxies but also out to high redshift (I. Cortzen et al., 2019; I. Shivaei & L. A. Boogaard, 2024; R. Chown et al., 2025; H. Shim et al., 2025). This prospect is particularly exciting because JWST is now routinely enabling high–resolution, high–sensitivity PAH imaging. Looking ahead, Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer satellite (SPHEREx; O. Doré et al., 2018) will deliver full–sky spectral PAH maps, opening the door to uniform cold ISM mapping across all nearby galaxies. However, for these applications to be reliable, we need to identify where the relation between PAH and cold gas breaks down, beyond classical H II regions, for example in AGN-dominated centers and zones of strong shear or shocks.
In this Letter, we examine regions that deviates significantly from the CO–7.7m PAH relation as calibrated by R. Chown et al. (2025) across the same PHANGS galaxies. We then assess the local environment, galactic dynamics, and excitation conditions of these outliers to isolate the physical processes that drive these deviations from the global correlation.
2 Observational Data
We focus on the same 70 galaxies from R. Chown et al. (2025) for which scaling relations between CO and 7.7m PAH emission were derived using high-resolution JWST imaging with the F770W filter and ALMA CO observations. These targets are from PHANGS–JWST Cycle 1 (GO 2107; PI: Lee; J. C. Lee et al., 2023; T. G. Williams et al., 2024) and Cycle 2 (GO 3707; PI: Leroy; R. Chown et al., 2025). Most of these galaxies lie on the star-forming main sequence, with stellar masses ranging from to 11.5 and at distances of Mpc.
2.1 JWST mid-IR imaging of PAH emission
For 19 galaxies from the PHANGS-JWST Cycle 1 program, we use JWST NIRCam and MIRI imaging data observed with F335M, F770W, F1000W, F1130W, and F2100W filters. The Cycle 1 data can be obtained using this link111https://archive.stsci.edu/hlsp/phangs or Williams, Thomas et al. (2023). The remaining 51 galaxies are from the PHANGS-JWST Cycle 2 program described in R. Chown et al. (2025), with limited coverage in filters. We use two mid-IR bands of F770W and F2100W.
The F335M, F770W and F1130W filters capture strong PAH emission features at 3.3, 7.7 and 11.3m. The F1000W (centered at 10m) is expected to be dominated by dust continuum (C. M. Whitcomb et al., 2023), though its morphology and intensity often track CO and H, similar to the PAH bands (A. K. Leroy et al., 2023). The F2100W at 21m captures dust continuum emission, including contributions from grains in thermal equilibrium and heated by the interstellar radiation field (B. T. Draine & A. Li, 2007; R. C. Kennicutt & N. J. Evans, 2012; A. K. Leroy et al., 2012; F. Belfiore et al., 2023a; H. Hassani et al., 2023; A. K. Leroy et al., 2023; D. Pathak et al., 2024). The resolution of F335M, F770W, F1000W, F1130W, F2100W imaging is 011, 027, 033, 038, and 067, respectively222Point spread functions for NIRCAM and MIRI, where the resolution of F2100W roughly translates into a physical scale of pc across the range of distances of our galaxy sample ( Mpc). Both of the Cycle 1 and Cycle 2 data have been reduced using the PHANGS-JWST pipeline (pjpipe333https://pjpipe.readthedocs.io/en/latest/). For comprehensive descriptions of the survey and the data reduction process, we refer readers to J. C. Lee et al. (2023), T. G. Williams et al. (2024), and R. Chown et al. (2025).
The F335M and F770W bands include contributions from the stellar continuum. For F335M, we use continuum-subtracted maps from H. Koziol et al. (in prep.), where the stellar continuum is estimated using F300M and F360M, based on the method of K. M. Sandstrom et al. (2023b). As for F770W, we used stellar continuum-subtracted F770W maps adopted in R. Chown et al. (2025), obtained by subtracting the flux in F200W (Cycle 1) or F300M (Cycle 2) scaled by a factor determined using the prescription of J. Sutter et al. (2024).
2.2 ALMA 12CO( imaging
We use 12CO() imaging from the PHANGS–ALMA survey, hereafter CO. Full descriptions of the sample, survey design, and image production process are provided in A. K. Leroy et al. (2021a) and A. K. Leroy et al. (2021b). The observations were conducted with the 12-m, 7-m, and total power arrays of ALMA. The CO maps have resolutions of 10-17, corresponding to physical scales of 25–150 pc for the galaxies in our sample. We use the first public release of the moment-0 maps, generated using an inclusive signal-masking scheme (“broad” masking; see A. K. Leroy et al., 2021b) designed to capture extended and faint emission and encompass all emission from each galaxy, at the native resolution. We also use maps of the CO effective line width () from the same public release, produced with the more restrictive “strict” mask, which includes only voxels highly likely to contain real emission.
2.3 ALMA 12CO( imaging
If outliers with abnormally high CO/PAH ratios are driven by CO excitation, the CO line ratios (12COCO) could increase. In order to test this, we use ALMA 12CO imaging from den Brok et al. (in preparation; 2022.1.01479.S) obtained using 12-m, 7-m, and total power antennas (see also J. den Brok et al., 2025). The observations were available for 14 of the 70 PHANGS galaxies. Among these 14 galaxies, 4 galaxies host significant outliers.
2.4 MUSE ionized gas emission
For PHANGS-JWST Cycle 1 galaxies, we leverage optical integral field spectroscopy from the PHANGS–MUSE survey (E. Emsellem et al., 2022) in order to look for signs of massive star formation. The PHANGS–MUSE survey observed all 19 of the JWST Cycle 1 galaxies with the Multi Unit Spectroscopic Explorer (MUSE) on the Very Large Telescope (VLT), covering the star-forming disks at a spatial resolution of 10 (20–100 pc) over 4800–9300 Å, enabling measurements of key emission lines including H and H. Emission line fluxes were derived by simultaneously fitting the simple stellar population models to continuum and a Gaussian distribution to gas emission lines. We use H maps from F. Belfiore et al. (2023b), which has been corrected for dust extinction using the Balmer decrement. As for the PHANGS-JWST Cycle 2 galaxies without MUSE coverage for the regions of interest, we adopt continuum-subtracted narrow-band H imaging from Razza et al. (in preparation).
2.5 Convolution and homogenization to a common pixel grid
To proceed with our analysis of comparing CO and PAH emission, we apply additional processing to the data. First, the higher-resolution JWST observations are convolved to match the coarser resolution of the CO maps. The convolution uses a kernel that transforms the JWST PSF into a circular Gaussian PSF, generated using the code jwst_kernels444https://github.com/francbelf/jwst_kernels, following the method of G. Aniano et al. (2011) and T. G. Williams et al. (2024). Then, the convolved images are re-gridded to align with the pixel grid of the CO data.
3 Outlier detection
We perform outlier detection using CO and stellar continuum subtracted F770W maps that has been resampled on a gird with one CO-beam spacing to reduce inter-sample covariance. We then identify pixels whose CO emission is anomalously high relative to the 7.7 m PAH emission. For each galaxy, we predict CO emission from the stellar–continuum–subtracted F770W emission () using the relation:
where the coefficients , , and are individual galaxy based fits from Table 4 of R. Chown et al. (2025). We then calculate residuals
Uncertainties on () are obtained by error propagation,
where and are per beam measurement errors, are the fit–parameter uncertainties, and is the intrinsic scatter of the relation from R. Chown et al. (2025).
To remain consistent with R. Chown et al. (2025), we limit the analysis to beam samples with . To avoid noise contamination and isolate significant outliers, we additionally require in both and . We then identify beam samples with and group adjacent beams into a contiguous region and compute inverse–variance–weighted statistics to calculate average residual and its uncertainty ( and ). We define a contiguous region as having abnormally high CO-to-PAH ratio if it satisfies the conservative detection threshold of . This results in a detection of outlier regions in galaxies out of 70 galaxies analyzed in R. Chown et al. (2025). If we relax the threshold to , additional outlier regions that are relatively small, composed of two to three CO beams, emerge in NGC 1512, NGC 1672, NGC 2835, NGC 4457, and NGC 4548. Throughout the paper, we adopt a more conservative threshold to focus on significant and obvious outliers.
Using the same procedure, we searched for low CO-to-PAH ratio outliers , which could signal CO-dark molecular or atomic gas (K. M. Sandstrom et al., 2023a; J. Kim et al., 2025). We did not identify such regions, likely because the CO data is too shallow and noisy to establish a significant deficit relative to the PAH emission.
In Figure 1 we show the measured ratio () for three galaxies, selected because their outlier regions are either spatially extended or located across diverse environments. The red contours in Figure 1 highlight regions where the observed CO flux exceeds the value expected from the 7.7m PAH emission by more than an order of magnitude. In Appendix A, we present the same figure for all 20 galaxies with such outliers, along with their properties, among the 70 galaxies included in our analysis. We note that most galaxies hosting outliers are classified as strongly or weakly barred (17 out of 20; R. J. Buta et al. 2015). However, outliers are not ubiquitous among barred systems, with only 35% of barred galaxies exhibiting outliers.
These abnormal regions represent significant outliers from the tight CO–7.7m PAH relation reported by R. Chown et al. (2025). They constitute 0.03-7% of the total CO-emitting area with an average of 1% (see Appendix A) and are spatially localized along bar lanes and centers without signs of massive star formation detected in H. NGC 1566 is the only galaxy where we find such regions at the edges of spiral arms. The preferential location of these regions within specific dynamical structures suggests that they are not noise-driven statistical outliers. Instead, the elevated CO-to-PAH ratios have a physical origin linked to local gas dynamics and/or radiation environments.
In order to investigate the origin of these high CO-to-PAH ratio regions, we define a “null test” control region with three requirements: 1) it lies near the outlier region within the same dynamical structure; 2) it contains only pixels with normal CO-to-PAH ratios (); and 3) it shows no obvious signs of recent star formation, so as to remain consistent with the outlier regions. Recent star formation is traced by extinction-corrected MUSE H for Cycle 1 galaxies and by narrowband H imaging for Cycle 2 galaxies. In addition to these, whenever possible, we also required the null test control region to have a CO flux comparable to that of the outlier region. The locations of these null test regions are shown in Appendix A.
In the following section, by comparing outlier regions to its null-test regions, we explore possible physical mechanisms, including suppressed PAH emission and enhanced CO emissivity that could produce such extreme deviations from the typical CO–PAH relation.
4 Origin of outliers
4.1 Suppressed PAH Emission Scenario
As described in Section 2.1, the JWST Cycle 1 data provide coverage across multiple filters, enabling measurements of key PAH emission features (at 3.3, 7.7, and 11.3m) as well as the dust continuum (at 10m and 21m). Here, we investigate whether the outlier regions identified in the CO-to-7.7m PAH ratio persist when other PAH or dust continuum dominated bands are used. We also compare PAH properties between the abnormal and null-test regions.
4.1.1 CO-to-mid-IR ratio using other PAH and dust continuum emission
Figure 2 shows the CO flux relative to that expected from other PAH- and dust continuum–dominated bands in NGC 4303, which is one of the galaxies with the most prominent CO-to-7.7m PAH anomalies. We estimate CO fluxes from 3.3 and 11.3m bands using the coefficients from R. Chown et al. (2025) ( and ), and from 10 and 21m bands using coefficients derived in this work following the same method ( and ).
The abnormal regions identified in CO-to-7.7m PAH ratio (red) consistently show elevated ratios across all PAH and continuum bands, with the largest excess observed at 3.3m and the smallest at 10m. Other galaxies hosting abnormal regions also show a similar behavior. The significant enhancement across multiple bands indicates that 7.7m is not particularly suppressed and the anomaly is more likely due to overluminous CO emission rather than a physical mechanism that suppresses multiple PAH bands and dust continuum emission at the same time. It is also interesting to note that shows the closest agreement with , even though m traces mostly PAH emission while m traces mostly dust continuum. Further supporting this picture, Appendix A shows three–color HST composites (F814W, F438W, F555W; J. C. Lee et al., 2022) in which the outlier regions display dust extinction similar to their surroundings (see also H. Faustino Vieira et al. 2025). This suggests that abnormal regions exhibit comparable dust content compared to their surroundings, making simultaneous suppression of PAH and dust emission an unlikely explanation for the elevated CO/PAH ratios.
4.1.2 Difference in PAH properties
PAH intensity ratios are sensitivity to molecule properties (e.g., size, asymmetry, composition, and charge state) and to local radiation field (S. Hony et al., 2001; F. Galliano et al., 2008; T. S. Y. Lai et al., 2020; A. Maragkoudakis et al., 2020). In particular, 11.3m/3.3m becomes lower as the spectral energy distribution of the illumination radiation becomes harder and PAH size distributions become smaller. The ratio of 11.3m/7.7m also decreases with harder radiation field and is very sensitive to changes of PAH charge variations (T. S. Y. Lai et al., 2022; J. Chastenet et al., 2023; B. T. Draine et al., 2021; B. S. Hensley & B. T. Draine, 2023b; D. Baron et al., 2024, 2025). Furthermore, the ratio of emission in PAH bands to the 21 m dust continuum is considered as a strong tracer of the fraction of dust mass in the form of PAH (J. Chastenet et al., 2019; J. Sutter et al., 2024; O. V. Egorov et al., 2023).
In Figure 3, we show median and 1 distribution of PAH band ratios to assess whether PAH properties differ between our abnormal and null-test regions across Cycle 1 galaxies. The fraction of dust mass in PAHs, , estimated from , shows no significant difference between abnormal and null-test regions. We note that the abnormal and null test regions also do not overlap with massive star-forming sites, where is typically suppressed due to PAH destruction by stellar feedback (J. Chastenet et al., 2023; O. V. Egorov et al., 2023, 2025; J. Sutter et al., 2024).
In the middle panel, we compare . While most galaxies follow a one-to-one relation, abnormal regions in NGC 1433 and NGC 4535 show somewhat lower compared to their null-test region, suggesting fragmentation into smaller grains or harder radiation field. In contrast, NGC 3627, shows a higher ratio in its abnormal region (see also (D. A. Dale et al., 2025)), possibly indicating PAH growth or coagulation onto dust grains, making 7.7m and 11.3m bands fainter (B. T. Draine et al., 2021; B. S. Hensley & B. T. Draine, 2023b). As most of these outlier regions lie along bar lanes with no signs of massive star formation, the differences in suggest that PAH size distributions may be altered by shocks or coagulation as gas and dust flow inward.
Lastly, we compare PAH ratios. The , often used as a tracer of PAH ionization (e.g., J. D. T. Smith et al., 2007; T. S. Y. Lai et al., 2022), has also been shown to correlate with (D. Baron et al., 2024). This implies that the majority of the variation in ratios is driven by the changing shape of the illuminating spectral energy distribution, where the ratio becomes higher as the radiation from old stellar population dominates (high and high ) compared to the radiation from young massive stars (low and low ). In NGC 1433 and NGC 4321, we find somewhat higher in abnormal regions ( dex), which may indicate an older stellar population dominating the radiation and/or more neutral PAHs, and therefore the lack of PAH emission compared to CO. However, in most galaxies, abnormal and null-test regions agree within 1.
While some abnormal regions show hints of altered PAH sizes or hardness of the radiation field, these trends are not consistent across systems. As the PAH ratios explored here are degenerate with multiple physical effects, additional spectroscopic tracers (e.g., H2 rotational and rovibrational lines, [Fe II] for shocks, and a wider suite of PAH features) will be essential to further investigate whether PAH emission is suppressed in these regions.
4.2 Elevated CO emissivity scenario
Almost all abnormal regions lie along the dust lanes in bars, sites where inflowing gas experiences streaming motions and strong shocks due to intersecting orbital families, causing abrupt changes in velocity (W. W. Roberts et al., 1979; E. Athanassoula, 1992; M. C. Sormani et al., 2015; J. A. Sellwood & A. Wilkinson, 1993). These dynamical processes can elevate the CO velocity dispersion and increase the integrated intensity either through line broadening or by reducing the optical depth, which boosts the CO emissivity (A. D. Bolatto et al., 2013; M. Gong et al., 2020; Y.-H. Teng et al., 2023). In addition, shocks and strong shear can raise the excitation temperature and collision rates, populating higher- levels and further enhancing the CO flux for a given molecular gas mass.
In this Section, to test if high CO-to-PAH ratio is due to elevated CO emissivity or excitation by comparing CO velocity dispersions and CO line ratios between outlier and null-test regions. If the high CO/PAH ratios are associated with higher velocity dispersions or elevated excitation temperature, it would indicate that the -to-PAH ratio between abnormal and null-test regions should remain similar and the CO–to– conversion factor () should be at least 10 times lower in abnormal regions compared to the regions with .
4.2.1 CO velocity dispersion
Centers of barred galaxies exhibit higher velocity dispersions compared to the galaxy disk or centers of non-barred galaxies (J. Sun et al., 2020). The elevated dispersions in barred environments reduce optical depth, raise the CO emissivity, and lower the (Y.-H. Teng et al., 2024). Using dust-based, spatially resolved measurements at 2 kpc scale for nearby PHANGS galaxies (I.-D. Chiang et al., 2024), Y.-H. Teng et al. (2024) found a strong anti-correlation between and velocity dispersion across both galaxy centers and disks, suggesting that the CO velocity dispersion directly traces changes in CO emissivity.
Figure 4 shows the CO luminosity-weighted average velocity dispersion () as a function of the median CO-to-PAH ratio () of each abnormal region and null-test region, across the 20 galaxies hosting CO/PAH outliers. On average, the abnormal regions show two times higher and times higher compared to those in null-test regions555When all pixels are considered and not just the null-test region, the distribution of strongly peaks at 1.. The bottom panel shows as a function of the median CO-to-PAH ratio (). We obtain using the Y.-H. Teng et al. (2024) prescription,
with measured within each region. We also compare the distributions of and , where . As abnormal regions are associated with higher velocity dispersions and thus lower , the large offset observed between the abnormal and null-test regions in distribution decreases when is used. We also show how high (or how low the ) should be in order to explain these outliers solely by enhanced CO-emissivity, under the assumption of constant -to-PAH ratio (dotted line) between abnormal and null-test regions. While the abnormal regions are associated with higher velocity dispersion, it appears it should be even higher (three times more) to explain these outliers from enhanced emissivity.
The bottom panel of Figure 4 shows CO velocity dispersion () as a function of molecular gas surface density () for three galaxies where the enhancement in CO velocity dispersion is prominent. We also divide the data into logarithmically spaced bins in and then compute median and 1 scatter of in each bin for pixels in abnormal null-test regions. Within the range of where the abnormal and null-test regions overlap, the abnormal regions generally exhibit higher , indicating higher CO emissivity that can lead to abnormally high CO-to-PAH ratios.
In Appendix B, the same figure for all 20 galaxies hosting outlier regions is shown. While in some galaxies the velocity dispersions at a given agree within , almost half of them show differences exceeding this level, particularly NGC 1566, NGC 3351, NGC 4321, NGC 2566, NGC 2997, NGC 4569, NGC 4579, NGC 6300. We also note that when prescription from (Y.-H. Teng et al., 2024) is used, between abnormal and null-test regions overlaps much more significantly than when a constant Galactic is adopted (triangles). This again supports the view where abnormally high CO-to-PAH ratios are driven by elevated CO emissivity. The abnormal and null-test regions contain comparable amounts of PAH emission for a given amount of .
4.2.2 CO excitation from multi-transition lines and CO isotopologues
High-resolution observations of CO multi-transitions and CO isotopologues observations would enable direct constraints on variations in CO line opacity and emissivity. For example the ratio between CO lines from and () increases in more highly excited gas, where higher kinetic temperature and density populate more efficiently than , enhancing CO() relative to CO(). However these lines are optically thick in molecular clouds and the ratio saturates.CO isotopologues, which are less opaque than 12CO, further help disentangle excitation from opacity effects. Their multi-transition line ratios provide sensitive probes of excitation, while 12CO/13CO ratios are sensitive to isotopologue abundance and/or optical depth.
Using CO isotopologues observations in their low rotational transitions, Y.-H. Teng et al. (2022) and Y.-H. Teng et al. (2023) have measured local variations in the barred-centers of NGC 3351, NGC 3627, and NGC 4321, which were found to be 4-15 times lower compared to the Galactic value. These three galaxies overlap with our sample that host abnormal regions, however, the current available data are not deep enough to detect 13CO in the regions of interest for NGC 3627 and NGC 4321. As for NGC 3351, the locations of high CO/PAH ratio is near the region with very low , located at where bar-driven inflow feeds the nucleus.
Using maps from J. den Brok et al. (2025, ; in preparation) for four galaxies hosting outliers (NGC 1300, NGC 1433, NGC 2566, and NGC 3627), we did not find significant differences in between our abnormal and null-test regions exceeding 1. Y.-H. Teng et al. (2023) have mapped 13CO()/() 12CO/13CO() in NGC 3351 including the outlier region. Figure 5 shows that outlier regions are associated with lower 13CO()/() and higher 12CO/13CO(). These trends suggest that high CO-to-PAH ratio is not due to higher excitation temperature and is consistent with reduced optical depth that allows more CO emission to escape, thereby boosting both the integrated CO intensity and the 12CO/13CO() ratio.
5 Summary
Under the conditions that dust and PAH grains are well mixed with gas and variations in the strength of the interstellar radiation field are modest, stochastically-heated PAH emission is expected to show correlation with the total amount of gas, often traced with CO in molecular gas-dominated galaxies. In line with this expectation, R. Chown et al. (2025) have found a tight and almost linear correlation between 7.7m PAH emission and CO flux on 100 pc scales, across 70 PHANGS-JWST galaxies.
In this work, we present localized departures from this CO–7.7m PAH relation, highlighting regions where the tight correlation breaks down. We have identified regions where the observed CO flux is more than 10 times brighter than that expected from 7.7m PAH emission. We find that such outliers are common, observed in 20 galaxies out of 70 explored in R. Chown et al. (2025). These regions with abnormally high are mostly located along bar lanes and in the centers of barred galaxies with an exception of NGC 1566 where we find such regions to exist in the concave part of the spiral arm as well. Interestingly, outlier regions do not show associated H emission, indicating that the abnormally high CO-to-PAH ratio is not due to massive stars dissociating PAHs (O. V. Egorov et al., 2023; J. Sutter et al., 2024).
We investigate the origin of the abnormally high CO-to-PAH ratios, which may arise from either suppressed 7.7m PAH emission or enhanced CO emissivity.
1. Suppressed 7.7m PAH emission scenario:
For the JWST Cycle 1 galaxies, we find that the abnormal regions, identified based on elevated CO-to-7.7m PAH ratios, also consistently exhibit elevated ratios when using other PAH (3.3 and 11.3m) and dust continuum bands (10 and 21m). The fact that the ratios are significantly enhanced across multiple PAHs and dust-continuum bands suggests that the abnormal CO-to-PAH ratios are more likely due to overluminous CO emission, rather than suppressed 7.7m PAH emission.
We also find that some abnormal regions either show bigger or smaller PAH size distributions or hints of PAHs being excited by softer radiation fields compared to their null-test regions. However, these trends are not consistent across the system.
2. High CO emissivity scenario:
Abnormal regions generally show higher CO velocity dispersions () compared to their null-test regions. Higher velocity dispersion leads to lower optical depth in the CO line, thereby increasing the CO emissivity for a given gas mass. This implies that the abnormal regions contains a similar amount of gas as their neighboring regions (as suggested by the similar PAH flux), but the elevated CO emissivity enhances the CO-to-PAH ratio.
Multi-transition CO and isotopologue data enables more direct constraints on opacity and excitation. While can rise with higher excitation, maps for four galaxies show no differences exceeding 1 between abnormal and null-test regions. In NGC 3351, isotopologue ratios show lower and higher in outlier regions, favoring reduced effective optical depth as the driver of the high CO/PAH instead of higher excitation temperature.
These results support the idea that variations in can occur on sub-kpc scales with changes of up to an order of magnitude, as observed in the Milky Way (L. Blitz et al., 1985) and centers of barred galaxies (Y.-H. Teng et al., 2022, 2023), and is not solely determined by large-scale galactic environments (e.g., bar versus disk) or metallicity.
In summary, we have explored two possible scenarios to explain the extreme outliers in CO-PAH relation. While the elevated CO velocity dispersion may suggest changes in CO emissivity as the driver of the enhanced CO-to-PAH ratios, this interpretation requires confirmation through deep observations of CO isotopologues and multi-transitions that can directly constrain variations in optical depth and excitation. Future JWST IFU observations targeting additional PAH bands and shock tracers, along with measurements of H2 lines, will be crucial to distinguish between these scenarios.
ADS/JAO.ALMA#2012.1.00650.S, ADS/JAO.ALMA#2013.1.00803.S, ADS/JAO.ALMA#2013.1.01161.S, ADS/JAO.ALMA#2015.1.00121.S, ADS/JAO.ALMA#2015.1.00782.S, ADS/JAO.ALMA#2015.1.00925.S, ADS/JAO.ALMA#2015.1.00956.S, ADS/JAO.ALMA#2016.1.00386.S, ADS/JAO.ALMA#2017.1.00392.S, ADS/JAO.ALMA#2017.1.00766.S, ADS/JAO.ALMA#2017.1.00886.L, ADS/JAO.ALMA#2018.1.01321.S, ADS/JAO.ALMA#2018.1.01651.S, ADS/JAO.ALMA#2018.A.00062.S, ADS/JAO.ALMA#2019.1.01235.S, ADS/JAO.ALMA#2019.2.00129.S, ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), NSC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
Appendix A Full list of galaxies hosting CO/PAH outliers
Table 1 lists physical properties of 20 galaxies hosting outlier regions, together with the fraction of outlier pixels relative to the entire CO-emitting area. From Figures 6 to 11, we show all the galaxies studied in R. Chown et al. (2025) that host regions with abnormally high CO-to-PAH ratios. When available, three-color composite HST images from the PHANGS–HST survey (J. C. Lee et al., 2022) are also included for some galaxies. Red contours highlight abnormal regions, whereas their null test regions are identified in black.
| (a) | (b) | (c) | (d) | (e) | (f) | (g) | ||
| Galaxy | JWST Cycle | Dist. | Incl. | Morphology | ||||
| [] | [] | [Gyr-1] | [Mpc] | [deg] | type | [%] | ||
| NGC1300 | Cycle 1 | 10.6 | 0.1 | -1.5 | 18.99 | 31.8 | Sbc | 3.03 |
| NGC1365 | Cycle 1 | 11.0 | 1.2 | -0.8 | 19.57 | 55.4 | Sb | 1.17 |
| NGC1433 | Cycle 1 | 10.9 | 0.1 | -1.8 | 18.63 | 28.6 | SBa | 4.03 |
| NGC1566 | Cycle 1 | 10.8 | 0.7 | -1.1 | 17.69 | 29.5 | SABb | 0.64 |
| NGC3351 | Cycle 1 | 10.4 | 0.1 | -1.2 | 9.96 | 45.1 | Sb | 0.25 |
| NGC3627 | Cycle 1 | 10.8 | 0.6 | -1.2 | 11.32 | 57.3 | Sb | 0.20 |
| NGC4303 | Cycle 1 | 10.5 | 0.7 | -0.8 | 16.99 | 23.5 | Sbc | 0.59 |
| NGC4321 | Cycle 1 | 10.7 | 0.6 | -1.2 | 15.21 | 38.5 | SABb | 0.06 |
| NGC4535 | Cycle 1 | 10.5 | 0.3 | -1.2 | 15.77 | 44.7 | Sc | 0.53 |
| NGC1097 | Cycle 2 | 10.8 | 0.7 | -1.1 | 13.58 | 48.6 | SBb | 2.04 |
| NGC2566 | Cycle 2 | 10.7 | 0.9 | -0.8 | 23.44 | 48.5 | Sb | 7.43 |
| NGC2903 | Cycle 2 | 10.6 | 0.5 | -1.1 | 10.0 | 66.8 | Sbc | 0.15 |
| NGC2997 | Cycle 2 | 10.7 | 0.6 | -1.1 | 14.06 | 33.0 | SABc | 0.04 |
| NGC3507 | Cycle 2 | 10.4 | -0.0 | -1.4 | 23.55 | 21.7 | SBb | 0.34 |
| NGC4569 | Cycle 2 | 10.8 | 0.1 | -1.7 | 15.76 | 70.0 | Sab | 0.70 |
| NGC4579 | Cycle 2 | 11.1 | 0.3 | -1.8 | 21.0 | 40.22 | Sb | 1.08 |
| NGC4731 | Cycle 2 | 9.5 | -0.2 | -0.7 | 13.28 | 64.0 | SBc | 1.04 |
| NGC4941 | Cycle 2 | 10.2 | -0.4 | -1.5 | 15.0 | 53.4 | SABa | 0.35 |
| NGC5643 | Cycle 2 | 10.3 | 0.4 | -0.9 | 12.68 | 29.9 | Sc | 1.18 |
| NGC6300 | Cycle 2 | 10.5 | 0.3 | -1.2 | 11.58 | 49.6 | SBb | 1.84 |
Note. — (a) (b), and (c) – Global stellar mass, star formation rate, and specific star formation rate from PHANGS–ALMA (A. K. Leroy et al., 2019, 2021a). (d) and (e) – Distance from G. S. Anand et al. (2021) and inclination from P. Lang et al. (2020). (f) – Galaxy morphology from G. Paturel et al. (1997). (g) – Fraction of outlier pixels relative to the entire CO-emitting region.
Appendix B CO velocity dispersions in individual galaxies
Figure 12 compares CO velocity dispersion () as a function of molecular gas surface density () between abnormal and null-test regions for all 20 galaxies hosting outlier regions.